OpenStudy (anonymous):

In a case pencils lie: 2 simple, 3 red, 2 blue, and 1 yellow. What probability of what the pencil taken out from a case will be colorful?

OpenStudy (anonymous):

@Hero

hero (hero):

I wonder what color are the two "simple" pencils.

OpenStudy (anonymous):

regular pencil when you writing in school not color ones

OpenStudy (anonymous):

most stuff you write with regular pencil

hero (hero):

Okay, so basically, we have to simplify this ratio to get the probability: \[\frac{\text{no of colored pencils}}{\text{total number of pencils}}\]

OpenStudy (anonymous):

\[\frac{ 6 }{ 8} = 7,5\]

OpenStudy (anonymous):

?

hero (hero):

You have the 6/8 part correct. I don't know what the 7,5 is for.

OpenStudy (anonymous):

i don't know what you answer?

hero (hero):

All you have to do is simplify 6/8 and you'll have the appropriate answer.

OpenStudy (anonymous):

can you type the answer i want see how it looks in mathematics in you version

OpenStudy (anonymous):

i got 6/8 what you got

hero (hero):

Do you know how to reduce fractions to simplest form?

OpenStudy (anonymous):

@Hero just say you version answer if you can with showing work

hero (hero):

\(\frac{6}{8}\) reduces to 3/4